Suppose you need $500 in 9 months. How much must you deposit today, if the deposit will earn interest at 6% per year, compounded monthly?
Answer(s): B
On the BAII Plus, press 9 N, 6 divide 12 = I/Y, 0 PMT, 500 FV, CPT PV. On the HP12C, press 9 n, 6 ENTER 12 divide i, 0 PMT, 500 FV, PV. Make sure the BAII Plus has the P/Y value set to 1.
What does the complement rule state?
Answer(s): A
The sum of probability must equal 1. So P(A) + P(not A) = 1.
Two events, A and B, are mutually exclusive if:
By definition, if A and B are mutually exclusive, they cannot occur simultaneously. Therefore, P(A and B) = 0 for mutually exclusive events A and B. Note that P(A and B) = P(A).P(B) if and only if A and B are independent.Clearly, if A and B are mutually exclusive, they cannot be independent unless at least one of the two has a zero probability of occurring (i.e. is an impossible event). Finally, the only way "P(A and B) = P(A) + P(B)" will be true is if P(A) = P(B) = 0. To see this, recall that P(A or B) = P(A) + P(B) - P(A and B). The only way for "P(A and B) = P(A) + P(B)" to hold would be to have P(A or B) = 0. This is possible only when P(A) = P(B) = 0.
Affirmative action commitments by industrial organizations have led to an increase in the number of women in executive positions. Satellite Office Systems has vacancies for two executives, which it will fill from among four women and six men. What is the probability that two woman are selected?
Answer(s): D
Probability of both positions being filled by women = 4/10*3/9 = 12/90.
Which one of the following is referred to as the population mean?
Answer(s): C
The normal symbol for the population mean is the Greek letter "Mu."
Post your Comments and Discuss CFA CFA I exam prep with other Community members:
Our website is free, but we have to fight against AI bots and content theft. We're sorry for the inconvenience caused by these security measures. You can access the rest of the CFA I content, but please register or login to continue.